OpenStudy (anonymous):

A line segment on a number line has its endpoints at -9 and 6. Find the coordinate of the midpoint of the segment. A. 1.5 * B. -1.5 C. 2 D. -3 Find the coordinates of the midpoint of HX given that H(-1,3) and X(7,-1). A. (3,1) * B. (0,4) C. (-3,1) D. (-4,0) please help!

OpenStudy (lncognlto):

For the second part of the question, the formula for the midpoint of a line is \[(\frac{ x1 + x2 }{ 2 }, \frac{ y1 + y2 }{ 2 })\]so \[\frac{ 7-1 }{ 2 }, \frac{ 3-1 }{ 2 }\], so that simplified will give the answer.

OpenStudy (lncognlto):

For the first part of the question,  Thus 9 + 6 will give us the full length of the line. 15 is then divided by 2. and the result is added to -9 to get the final answer.

OpenStudy (anonymous):

so is number two A?

OpenStudy (anonymous):

and my first is correct?

OpenStudy (lncognlto):

Yep, your first is correct and number two is A! :)